# Pragmatic Reasoning improves LLM Code Generation

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated impressive potential in translating natural language (NL) instructions into program code. However, user instructions often contain inherent ambiguities, making it challenging for LLMs to generate code that accurately reflects the user’s true intent. To address this challenge, researchers have proposed to produce multiple candidates of the program code and then rerank them to identify the best solution. In this paper, we propose CodeRSA, a novel code candidate reranking mechanism built upon the Rational Speech Act (RSA) framework, designed to guide LLMs toward more comprehensive pragmatic reasoning about user intent. We evaluate CodeRSA using one of the latest LLMs on a popular code generation dataset. Our experiment results show that CodeRSA consistently outperforms common baselines, surpasses the state-of-the-art approach in most cases, and demonstrates robust overall performance. These findings underscore the effectiveness of integrating pragmatic reasoning into code candidate reranking, offering a promising direction for enhancing code generation quality in LLMs.
@@ 1 Introduction
Recent advances in generative large language models (LLMs) have demonstrated their impressive ability to generate program code from user-provided natural language instructions (Liu et al., 2024b; Coignion et al., 2024). However, given the intrinsic complexities of coding and the potential ambiguities in user input, producing code in a single attempt may fail to explore the vast solution space, overlooking correct or higher-quality solutions (Liu et al., 2024a). A standard practice to address this shortcoming is to sample multiple solutions, which we refer to as “code candidates” (Chen et al., 2021; Brown et al., 2024), and to rerank them. Researchers have proposed various reranking strategies for code candidates, broadly divided into execution-driven and content-driven approaches. Due to the safety-risks associated with execution-driven approaches (Yetiştiren et al., 2023), we here focus on content-driven methods. These evaluate the generated text, often relying on token-level probabilities. For example, Coder reranking scores each candidate based on the cumulative probability of its tokens, sometimes however favoring “degenerate solutions” (generic or repetitive code) with disproportionately high token probabilities (Zhang et al., 2023).
Figure 1: A comparison of our approach CodeRSA (top) compared to the CoderReviewer method (bottom).
@@
When viewing code generation as a communicative process in which an LLM listens to the user’s intentions (Ouyang et al., 2022), Coder reranking evaluates candidate solutions solely from the listener’s perspective. Yet, research on human communication suggests that effective listeners reason about the speaker (who in turn reasons about the listener) Grice (1975). Frank and Goodman (2012; 2016) provided a principled method for quantifying this process based on a probabilistic framework based on game-theoretic notions, called the Rational Speech Act (RSA) framework. Pu et al. (2020, 2024) demonstrated the RSA framework’s effectiveness on program generation for a simple domain, while Schuster et al. (2024) reported negative results on a spreadsheet domain. One aspect that has held back RSA models from scaling up to realistic use cases is the computational overhead (Pu et al., 2024): it requires reasoning about the set of alternative instructions that the speaker could have given and about the set of alternative pieces of code that could solve the problem, which is very computationally expensive. Zhang et al. (2023) therefore proposed CoderReviewer reranking as a simplified scalable approach that simplifies these probability estimation processes over alternatives. However, it comes at the cost of not fully modelling the dialogic, interactive reasoning that can emerge when speaker and listener exchange information.
@@
This work proposes CodeRSA, enabling LLMs to reason as pragmatic listeners and rank code candidates based on the user’s underlying intentions. It addresses the probability estimates for the set of alternative code candidates and alternative utterances via a sampling approach. CodeRSA generates multiple code candidates, and then generates additional instructions for each candidate, forming a set of potential instructions (including the original one), as illustrated in Fig. 1: the literal listener $L_{0}$ estimates the probability of each code candidate given each instruction from the potential instruction set. The pragmatic speaker $S_{1}$ then applies a special normalization on these probabilities to quantify how specifically an instruction fits the generated code. By comparing these pragmatic speaker scores for the original instruction across all candidates, the pragmatic listener identifies the code candidate that most closely matches the original instruction from the speaker’s perspective, finalizing the reranking process (see Fig. 1).
@@
We conducted experiments using CodeRSA with Llama-3-8B-Instruct, one of the latest language models from the Llama family (Grattafiori et al., 2024) on OpenAI’s HumanEval benchmark (Chen et al., 2021). Our experiment results reveal that CodeRSA reliably outperforms the Coder and CoderReviewer reranking methods. Our qualitative analysis illustrates how the CodeRSA reranking enables better candidate selection, promoting a more comprehensive understanding of user intent.
@@ 2 Related Work
Natural Language to Code. Previous research has extensively explored generating code from natural language using neural network models (Ling et al., 2016; Rabinovich et al., 2017; Hayati et al., 2018). Recently, large language models (LLMs) have propelled significant advances in this area, driven by the transformer Vaswani (2017) architecture and large-scale pretraining. Their performance on code generation tasks often surpasses that of traditional models, and in many cases even rivals human programmers (Ni et al., 2024; Becker et al., 2023). Moreover, a recent study shows that LLMs also exhibit strong performance in code summarization, effectively translating code snippets into text (Akib et al., 2024).
Diversity Sampling in Code Generation. In a prior study, Chen et al. (2021) found that allowing the model to generate more candidates significantly raises the probability of yielding a correct answer. This practice, which encourages the model to produce a broader range of potential outputs, is often referred to as diversity sampling.
Code Reranking Methods. When LLMs generate multiple code candidates in response to the user instruction, execution-driven reranking methods such as CodeT (Chen et al., 2022) and AgentCoder (Huang et al., 2024) involve running test cases to gauge each candidate’s correctness. Although often effective, execution-driven approach can introduce additional safety risks and practical obstacles, such as the potential for malicious code execution, environment setup overhead, and resource constraints (Yetiştiren et al., 2023; Khoury et al., 2023). In contrast, content-driven reranking methods are far more versatile because they do not rely on execution and are not even confined to coding tasks.
Coder reranking. In prior work, Chen et al. (2021) reranked code candidates by estimating $\text{P}(c\mid i)$, where $c$ denotes the generated code candidate and $i$ denotes the given instruction. This process can also be called Coder reranking because the LLM is a mere Coder that estimates the candidate probability based on the corresponding instruction.
When using an LLM to estimate conditional probabilities, we compute the probability of each token iteratively. For example, in Coder reranking, the model processes candidate’s tokens from left to right: at each step, it calculates the probability of the current token given the instruction and the previously generated tokens, then appends that token to the context before moving on. The product of these sequential probabilities across all tokens yields the overall probability of the code candidate under the given instruction:
@@
Zhang et al. (2023) added the concept of a reviewer to Coder reranking (i.e., reassessing whether the instruction matches the generated code candidate). Formally, CoderReviewer metric is represented as follows:
CoderReviewer $=\text{P}(c\mid i)\cdot\text{P}(i\mid c)$
@@
The prompt positions of the instruction and code candidates are reversed in the reviewer’s implementation. Thus, the code-generation task is reformulated as an instruction-generation task to calculate the probability of a given instruction. The CoderReviewer metric is also considered a specialized form of maximum mutual information (Li and Jurafsky, 2016), measuring the bidirectional alignment between the generated code and the input instruction.
@@ 3 CodeRSA
This section introduces CodeRSA, an approach that employs the Rational Speech Act (RSA) framework to enhance the reranking of candidate code snippets. CodeRSA extends the models proposed by Cohn-Gordon et al. (2019) and Schuster et al. (2024). The core innovation in CodeRSA arises from the pragmatic listener, which is responsible for selecting and reranking code candidates. It does so by imagining how a “pragmatic speaker” would choose an instruction that best distinguishes the intended code among various potential instructions.
A pragmatic speaker evaluates how effectively a potential instruction communicates the intended code candidate. Within the RSA framework, this evaluation process necessitates an explicit calculation of probabilities. However, given the infinite combinations of possible instructions and code candidates, developing a robust and comprehensive probability estimation strategy becomes essential. To address this challenge, CodeRSA generates additional instructions for each candidate, thereby constructing an expanded set of potential instructions (including the original). Finally, by considering the potential instruction set, a pragmatic speaker can select the most appropriate instruction for each code candidate based on the motivation to effectively convey the message to a literal listener.
At the foundation of CodeRSA is the literal listener, who estimates the probability of each code candidate by interpreting the user’s instruction word for word, without inferring deeper speaker intent.
Literal Listener. A literal listener (denoted $L_{0}$) represents the simplest level of reasoning in the RSA framework. It interprets utterances solely according to their literal meaning, without any higher-level pragmatic reasoning. Theoretically, let $c$ be a candidate code and $i$ be an instruction, then:
$\text{P}_{L_{0}}(c\mid i)\propto\llbracket i\rrbracket(c),$
where $\text{P}_{L_{0}}(c\mid i)$ denotes the literal listener’s estimation of probability of candidate $c$ given instruction $i$. The interpretation function $\llbracket i\rrbracket$ maps an instruction $i$ to a function that takes a code candidate $c$ and returns 1 if $c$ correctly implements $i$, and 0 otherwise.
Likewise, most LLMs generate code by conditioning only on the user’s input instruction, which motivates CodeRSA’s use of an LLM to instantiate this literal listener:
@@
The literal listener provides CodeRSA’s baseline interpretation of user instructions, serving as the foundation for all subsequent reasoning. Note that the Coder reranking can also be considered a literal listener. For convenience, we refer to the logarithm of $\text{P}_{L_{0}}(c\mid i)$ as the Coder score.
@@
Pragmatic Speaker. In the RSA framework, the pragmatic speaker (denoted $S_{1}$) is primarily responsible for determining whether an instruction $i$ effectively conveys the intended meaning of a candidate code $c$ to the literal listener. Formally, a pragmatic speaker can be defined as:
$\text{P}_{S_{1}}(i\mid c)=\frac{\exp\left(\log\text{P}_{L_{0}}(c\mid i)-C(i)\right)}{\sum_{i^{\prime}}\exp\left(\log\text{P}_{L_{0}}\left(c\mid i^{\prime}\right)-C\left(i^{\prime}\right)\right)},$
Here, $\operatorname{C}(i)$ denotes a cost function that quantifies the “expense” of employing a particular instruction. This formula quantifies the pragmatic speaker’s estimation of the probability that a particular instruction will be used to describe a code candidate, as interpreted from the literal listener’s perspective. Note that an idealized RSA approach would consider every possible instruction $i^{\prime}$, which is infeasible in practice. Instead, CodeRSA leverages the fact that ambiguities in the original instruction often surface in the code candidates generated by LLMs, thereby enabling a sample-based estimation.
Given a user-provided instruction $i_{0}$, we request multiple candidate solutions $\{c_{1},\dots,c_{n}\}$ from the LLM. Each candidate $c_{j}$ may capture a slightly different interpretation of $i_{0}$. We then derive one or more instructions from each candidate $c_{j}$, forming a set of potential instructions $\{i_{0},i_{1},\ldots,i_{m}\}$, which we denote by $I$. This approach ensures that we sample relevant alternative instructions directly tied to how the model interprets the original instruction. Potential instructions then constitute a finite sample set that approximates the otherwise infinite instruction space for pragmatic speaker estimation.
@@
To simplify the model and focus on core pragmatic reasoning, we assume a uniform cost for all instructions, which effectively cancels out during normalization. However, a more detailed modeling of the cost function may provide additional insights, a point we further discuss in the Section 6. A pragmatic speaker then can be defined in a simplified form as:
$\text{P}_{S_{1}}(i\mid c)=\frac{\text{P}_{L_{0}}(c\mid i)}{\sum_{i^{\prime}\in I}\text{P}_{L_{0}}(c\mid i^{\prime})},$
In practice, the distribution of $\text{P}_{L_{0}}(c\mid i)$ is often highly skewed, with a few high probability candidates dominating and a large spread in probability estimates. To address this, we here propose to instead work with log probabilities, which in practice brings low probability alternatives closer together and gives more weight to the relative trends among instruction–code pairs. The pragmatic speaker’s preference score for a given instruction is hence quantified as:
$\text{R}_{S_{1}}(i\mid c)=-\frac{\log\text{P}_{L_{0}}(c\mid i)}{\sum_{i^{\prime}\in I}\log\text{P}_{L_{0}}(c\mid i^{\prime})}\;,$
$\text{R}_{S_{1}}(i\mid c)\;\in\;(-1,\,0).$
In log space, after normalizing with respect to log probabilities, smallest numbers will now correspond to the events with the highest probability. Therefore, we invert the sign to negative in order to still choose maximal score instructions. Notably, if $c$ is overly generic or appears plausible under multiple instructions, then the denominator $\sum_{i^{\prime}\in I}\log\text{P}_{L_{0}}(c\mid i^{\prime})$ increases, resulting in a lower overall score. This log-based metric prevents inflated scores for candidates that might superficially fit many different instructions, thereby ensuring that each instruction is evaluated on a relatively fair basis. An ablation study (Appendix A.3) confirmed that omitting the logarithmic transformation significantly degrades normalization and impairs CodeRSA’s reranking capability, with Appendix A.4 offering geometric mean–based theoretical support.
Overall, at the pragmatic speaker level, CodeRSA leverages the code candidates themselves to sample potential instructions, capturing the diversity of possible interpretations. It then uses the literal listener’s estimation to quantify how effectively each potential instruction reflects the meaning of a code candidate.
@@
Pragmatic Listener. The pragmatic listener (denoted $L_{1}$) re-examines the original instruction $i_{0}$ across all candidates, completing the backward reasoning guided by the pragmatic speaker’s preferences. In prior work, Degen (2023) defined a pragmatic listener as:
@@
Here, $\text{P}(c)$ denotes the prior probability of a given code candidate. We make the simplifying assumption that all candidates are equally likely a priori, which allows us to treat $\text{P}(c)$ as a constant and omit it from the calculations. Note that this uniform prior assumption may not hold in real-world scenarios, as some code might be inherently more common than others. We will discuss this assumption further in Section 6. Consequently, in CodeRSA, the pragmatic listener reranks candidates by their pragmatic speaker scores:
$\text{P}_{L_{1}}(c\mid i)\propto\text{R}_{S_{1}}(i\mid c),$
so that the candidate with the highest pragmatic speaker score for $i_{0}$ is selected as optimal.
From CodeRSA’s reasoning process, it follows that, for each candidate, the LLM is invoked $m+1$ times, where $m$ represents the number of generated instructions. Consequently, the overall computational complexity of CodeRSA is $\mathcal{O}\bigl{(}n(m+1)\bigr{)}$, which grows quadratically when both $n$ and $m$ increase equally.

### v3

@@ Abstract
Large Language Models (LLMs) have demonstrated impressive potential in translating natural language (NL) instructions into program code. However, user instructions often contain inherent ambiguities, making it challenging for LLMs to generate code that accurately reflects the user’s true intent. To address this challenge, researchers have proposed approaches that produce multiple candidates of the program code and then rerank them to identify the best solution. In this paper, we propose CodeRSA, a novel code candidate reranking mechanism built upon the Rational Speech Act (RSA) framework, designed to guide LLMs toward more comprehensive pragmatic reasoning about user intent. We evaluate CodeRSA using Llama-3-8B-Instruct and Qwen-2.5-7B-Instruct on two widely used code generation benchmarks, HumanEval and MBPP. Our experiment results show that CodeRSA consistently outperforms common baselines, surpasses the state-of-the-art approach in most cases, and demonstrates robust overall performance. These findings underscore the effectiveness of integrating pragmatic reasoning into code candidate reranking, offering a promising direction for enhancing code generation quality in LLMs.
@@ 1 Introduction
Recent advances in generative large language models (LLMs) have demonstrated their impressive ability to generate program code from user-provided natural language instructions (Liu et al., 2024b; Coignion et al., 2024). However, given the intrinsic complexities of coding and the potential ambiguities in user input, producing code in a single attempt may fail to explore the vast solution space, overlooking correct or higher-quality solutions (Liu et al., 2024a). A standard practice to address this shortcoming is to sample multiple solutions, which we refer to as code candidates (Chen et al., 2021; Brown et al., 2024), and to rerank them. Researchers have proposed various reranking strategies for code candidates, broadly divided into execution-driven and content-driven approaches. Due to the unreliability of automatically generated test suites (Chen et al., 2022) and the potential safety risks associated with code execution (Yetiştiren et al., 2023), we focus instead on content-driven methods, which evaluate the generated text, often relying on token-level probabilities. For example, Coder Reranking scores each candidate based on the cumulative probability of its tokens, sometimes however favoring ‘degenerate solutions” (generic or repetitive code) with disproportionately high token probabilities (Zhang et al., 2023a).
Figure 1: A comparison of our approach CodeRSA (top) compared to CoderReviewer (bottom).
@@
When viewing code generation as a communicative process in which an LLM listens to the user’s intentions (Ouyang et al., 2022), Coder Reranking evaluates candidate solutions solely from the listener’s perspective. Yet, research on human communication suggests that effective listeners reason about the speaker (who in turn reasons about the listener) Grice (1975). Frank and Goodman (2012; 2016) provided a principled method for quantifying this process based on a probabilistic framework rooted in game-theoretic notions, called the Rational Speech Act (RSA) framework. Pu et al. (2020, 2024) demonstrated the effectiveness of the RSA framework for program generation in a simple domain (regular-expression synthesis), while Schuster et al. (2024) reported negative results on a spreadsheet domain. One aspect that has held back RSA models from scaling up to realistic use cases is the computational overhead (Pu et al., 2024): It requires reasoning about a whole set of alternative instructions that the speaker could have given and about the set of alternative pieces of code that could solve the problem, which is computationally expensive. Zhang et al. (2023a) therefore proposed CoderReviewer Reranking as a scalable approach that simplifies these probability estimation processes over alternatives. However, this comes at the cost of not fully modelling the dialogic, interactive reasoning that can emerge when speaker and listener exchange information.
@@
In this paper, we propose CodeRSA, enabling LLMs to reason as pragmatic listeners and rank code candidates based on the user’s underlying intentions. It addresses the probability estimates for the set of alternative code candidates and alternative utterances via a sampling approach. CodeRSA generates multiple code candidates, and then generates additional instructions for each candidate, forming a set of potential instructions (including the original one), as illustrated in Fig. 1.
Following the RSA framework, the literal listener $L_{0}$ first estimates the probability of each code candidate given each potential instruction. The pragmatic speaker $S_{1}$ then normalizes these probabilities to measure how specifically an instruction fits the generated code. Finally, by comparing these pragmatic speaker scores for the original instruction across all candidates, the pragmatic listener identifies the code candidate that best aligns with the user’s intent, completing the reranking process (see Fig. 1).
A challenge arises when many instructions are semantically equivalent but differ only in surface form. Applying RSA directly in such cases can lead to an overinterpretation of the formulation choice: The reasoning process is forced to treat near-identical descriptions as distinct alternatives, which were chosen for a reason of differentiating from other meaning alternatives. This fragments probability mass and reduces accuracy. To mitigate this, CodeRSA employs a clustering step. It groups semantically equivalent descriptions using an LLM-based equivalence test, ensuring that pragmatic reasoning emphasizes genuine differences in meaning rather than superficial wording (see Fig. 1, Pragmatic Speaker Part, where the main cluster refers to the one containing the original description $I_{0}$).
@@
We conducted experiments using CodeRSA with Llama-3-8B-Instruct, one of the latest language models from the Llama family (Grattafiori et al., 2024), and Qwen-2.5-7B-Instruct, a recent instruction-tuned model from the Qwen series (Yang et al., 2024), on two widely used code generation benchmarks: OpenAI’s HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021). Our experimental results demonstrate that CodeRSA reliably outperforms the Coder and CoderReviewer Reranking methods. Our qualitative analysis reveals how the CodeRSA enables better candidate selection, promoting a more comprehensive understanding of user intent.
@@ 2 Related Work
Natural Language to Code. Previous research has extensively explored generating code from natural language using neural network models (Ling et al., 2016; Rabinovich et al., 2017; Hayati et al., 2018). Recently, large language models (LLMs) have propelled significant advances in this area, driven by the transformer Vaswani (2017) architecture and large-scale pretraining. Their performance on code generation tasks often surpasses that of traditional models, and in many cases even rivals human programmers (Ni et al., 2024; Becker et al., 2023). A recent study shows that LLMs also exhibit strong performance in code summarization, effectively translating code snippets into text (Akib et al., 2024).
Code Reranking Methods. Execution-driven reranking methods such as CodeT (Chen et al., 2022) and AgentCoder (Huang et al., 2024) evaluate code candidates by running them against automatically generated test suites. Although often effective, these methods rely on the availability and reliability of test suites, which are frequently incomplete or difficult to construct, and executing untrusted code can pose safety risks (Yetiştiren et al., 2023; Khoury et al., 2023). In contrast, content-driven reranking methods are far more versatile because they do not rely on execution and are not even confined to coding tasks.
Coder Reranking. Chen et al. (2021) rerank code candidates by estimating $\text{P}(c\mid i)$, where $c$ denotes the generated code candidate and $i$ denotes the given instruction. This process can also be called Coder Reranking because the LLM is a mere Coder that estimates the candidate probability based on the corresponding instruction. When using an LLM to estimate conditional probabilities, we compute the probability of each token iteratively. For example, in Coder Reranking, the model processes a candidate’s tokens from left to right: At each step, it calculates the probability of the current token given the instruction and the previously generated tokens, then appends that token to the context before moving on. The product of these sequential probabilities across all tokens yields the overall probability of the code candidate under the given instruction:
@@
Zhang et al. (2023a) introduced the idea of augmenting Coder Reranking with a reviewer, which jointly considers how likely a code candidate is under the instruction and how well the instruction is supported by the code. Formally, the CoderReviewer conditional probability is defined as:
$\text{P}_{\text{CR}}(c\mid i)~\propto~\text{P}_{\text{LLM}}(c\mid i)\cdot\text{P}_{\text{LLM}}(i\mid c)$
@@
By switching the positions of the instruction and code in the conditional formulation, the second term can be interpreted as reformulating the code-generation task as an instruction-generation task. This bidirectional formulation can be viewed as a specialized form of maximum mutual information (Li and Jurafsky, 2016).
@@ 3 CodeRSA
In this section, we introduce CodeRSA, an approach that builds on the Rational Speech Act (RSA) framework to enhance the reranking of candidate code snippets. CodeRSA extends the models proposed by Cohn-Gordon et al. (2019) and Schuster et al. (2024). The core innovation in CodeRSA arises from the pragmatic listener, which is responsible for selecting and reranking code candidates. It does so by imagining how a pragmatic speaker would choose an instruction that best distinguishes the intended code among various potential instructions.
Literal Listener. A literal listener (denoted $L_{0}$) represents the simplest level of reasoning in the RSA framework. It interprets utterances solely according to their literal meaning, without any higher-level pragmatic inference. Let $c$ denote a candidate program and $i$ a user instruction. Then:
@@
where $\text{P}_{\text{LLM}}(c\mid i)$ is the probability assigned by the LLM to candidate $c$ given instruction $i$. In an idealized RSA setting, the literal listener would evaluate all possible programs, but since the space of programs is unbounded, we approximate it by sampling a finite set of candidate codes from the LLM.
We additionally define a candidate prior distribution obtained by querying the LLM without any instruction context:
$\text{P}_{\text{prior}}(c)=\text{P}_{\text{LLM}}(c\mid\varnothing).$
This prior reflects how plausible a candidate program is in general, independent of the specific user instruction.
@@
Pragmatic Speaker. In the RSA framework, the pragmatic speaker (denoted $S_{1}$) is primarily responsible for determining whether an instruction $i$ effectively conveys the intended meaning of a candidate $c$ to the literal listener. Formally, a pragmatic speaker can be defined as:
$\text{P}_{S_{1}}(i\mid c)=\frac{\exp\left(\log\text{P}_{L_{0}}(c\mid i)-C(i)\right)}{\sum_{i^{\prime}}\exp\left(\log\text{P}_{L_{0}}\left(c\mid i^{\prime}\right)-C\left(i^{\prime}\right)\right)}.$
Here, $C(i)$ denotes a cost function for using instruction $i$. In an ideal RSA setting, the normalization spans every possible instruction $i^{\prime}$, which is intractable for code generation. To approximate this space in practice, we take the sampled candidate codes as anchors and derive $m$ alternative instructions from each of the $n$ code candidates, together with the original instruction $i_{0}$, yielding a finite task-relevant instruction set $I=\{i_{0},i_{1},\ldots,i_{mn}\}$. This construction provides a principled approximation of the otherwise infinite instruction space while keeping RSA’s normalization meaningful.
@@
To simplify the model and focus on core pragmatic reasoning, we assume a uniform cost for all instructions, which effectively cancels out during normalization. A detailed modeling of the cost function may provide additional insights, a point we further discuss in Section 6. A pragmatic speaker then can be defined in a simplified form as:
$\text{P}_{S_{1}}(i\mid c)=\frac{\text{P}_{L_{0}}(c\mid i)}{\sum_{i^{\prime}\in I}\text{P}_{L_{0}}(c\mid i^{\prime})}.$
@@
Pragmatic Listener. The pragmatic listener (denoted $L_{1}$) re-examines the original instruction $i_{0}$ across all candidates, completing the backward reasoning guided by the pragmatic speaker’s preferences. In the standard RSA formulation (Degen, 2023), a pragmatic listener is defined as:
@@
where $\text{P}(c)$ denotes the prior probability of candidate $c$.
In practice, directly multiplying a normalized distribution by $\text{P}(c)$ can distort the allocation of probability mass, as the prior may dominate post hoc. Instead, CodeRSA incorporates priors via a candidate-specific temperature applied before normalization at the speaker stage. Let $z_{c}$ be the within-task standardized log prior of candidate $c$ (estimated from the LLM without conditioning context), and define a candidate-specific temperature as
$\tau_{c}\;=\;e^{-\alpha z_{c}},\qquad\alpha\geq 0,\quad\tau_{c}>0,$
where $\alpha$ controls how strongly the prior influences the temperature scaling. A higher prior (larger $z_{c}$) yields a smaller temperature ($\tau_{c}<1$) and thus a sharper distribution over alternatives. Candidates with higher priors therefore emphasize their most confident clusters more strongly, typically those that align best with the original instruction (e.g., the “main cluster”), giving them a comparative advantage during reranking.
With this calibration, the pragmatic speaker used by the listener is
$\text{P}_{S_{1}}(i\mid c;\tau_{c})\;=\;\frac{\bigl(\text{P}_{L_{0}}(c\mid i)\bigr)^{1/\tau_{c}}}{\sum_{i^{\prime}\in I}\bigl(\text{P}_{L_{0}}(c\mid i^{\prime})\bigr)^{1/\tau_{c}}},$
which reduces to the standard RSA speaker when $\alpha=0$ (thus $\tau_{c}=1$). Finally, the pragmatic listener ranks candidates with respect to the original instruction:
$\text{P}_{L_{1}}(c\mid i_{0})\;\propto\;\text{P}_{S_{1}}(i_{0}\mid c;\tau_{c}).$
This formulation preserves the spirit of RSA while integrating priors in a stable and interpretable manner: Rather than post-hoc reweighting, priors act as adaptive temperatures that shape the pragmatic reasoning process upstream of normalization. In our experiments, we treat $\alpha$ as a tunable hyperparameter and find that performance is stable across a broad range of values (see Section 5).
Clustering Paraphrases. While the basic RSA formulation operates directly over the instruction set $I$, it can suffer from over-interpreting superficial variations in wording when many instructions are semantically equivalent and differ only in surface form. In such cases, RSA allocates probability mass across paraphrases as if they were meaningful distinct alternatives, diluting the signal and reducing accuracy.
Semantically equivalent instructions:
 “return the sum of a list of integers”
 “compute the total of all integers in a list”
Non-equivalent instruction:
 “return the product of a list of integers”
Table 1: Examples of equivalent and non-equivalent instructions from MBPP. The first group expresses the same semantics, while the second differs in meaning.
To mitigate this, CodeRSA employs a semantic clustering. Candidate instructions are grouped into semantic clusters $\mathcal{C}=\{C_{1},\ldots,C_{K}\}$ using an LLM-based equivalence test (implementation details in Section 4), so that pragmatic reasoning operates over clusters rather than individual instructions. This ensures that comparisons emphasize genuine differences in meaning rather than superficial variation.
For a candidate $c$ and cluster $C_{k}$, the literal listener probability is aggregated as:
$\text{P}_{L_{0}}(c\mid C_{k})=\begin{cases}\text{P}_{L_{0}}(c\mid i_{0}),&\text{if }i_{0}\in C_{k},\\[4.0pt] \frac{1}{|C_{k}|}\!\sum_{i\in C_{k}}\!\text{P}_{L_{0}}(c\mid i),&\text{otherwise.}\end{cases}$
The pragmatic speaker distribution over clusters then becomes:
$\text{P}_{S_{1}}(C_{k}\mid c;\tau_{c})\;=\;\frac{\bigl(\text{P}_{L_{0}}(c\mid C_{k})\bigr)^{1/\tau_{c}}}{\sum_{C_{k^{\prime}}\in\mathcal{C}}\bigl(\text{P}_{L_{0}}(c\mid C_{k^{\prime}})\bigr)^{1/\tau_{c}}},$
where the candidate-specific temperature $\tau_{c}=e^{-\alpha z_{c}}$ incorporates priors.
Finally, the pragmatic listener reranks candidates with respect to the cluster $C^{\ast}$ containing the original instruction $i_{0}$:
$\text{P}_{L_{1}}(c\mid i_{0})\;\propto\;\text{P}_{S_{1}}(C^{\ast}\mid c;\tau_{c}).$
This extension preserves the primacy of the original instruction while preventing RSA from over-differentiating among paraphrases. Moreover, the integration of priors through adaptive temperatures ensures that the reranking remains calibrated against candidate plausibility.